OpenStudy (anonymous):

How would I solve this?: f(x+1) =4x^2+3x-2

OpenStudy (littlebird):

Im guessing.... where ever x is, replace it with (x+1) Simplify

OpenStudy (triciaal):

welcome to Openstudy!

OpenStudy (freckles):

are you wanting to find f(x)?

OpenStudy (jeansweets):

are you asking for 4x squared?

OpenStudy (triciaal):

to solve find the values of x that makes this = 0 or given f(x) = 4x^2 + 3x -2 find f(x+1) ?

OpenStudy (triciaal):

please post the complete question.

OpenStudy (freckles):

\[f(x+1)=g(x) \\ \text{ if you want \to find } f(x) \text{ given something like } f(x+1)=g(x) \\ \text{ then replace } x+1 \text{ with } u \\ \text{ so you have } u=x+1 \\ \text{ subtract 1 on both sides } \\ u-1=x \\ \text{ so you ahve } f(x+1)=g(x) \text{ gives } f(u)=g(u-1) \\ \text{ and then evaluate } f(u)=g(u-1) \text{ at } x \\ \text{ and you have } \\ f(x)=g(x-1)\] g(x) represents the right hand side of you equality by the way